OpenStudy (anonymous):

How many possible pairs of prime numbers are there if the difference between the two prime numbers is 2 and both numbers are less than 50? list all the pairs.

myininaya (myininaya):

so a prime is odd except for 2 let's assume the prime number is greater than 2 since there is no prime number such that the different of the prime number and 2 is 2 ok we have some odd number in the form of n=2k+1 and another prime m=2j+1 assume 50>m>n>0 (assuming we are talking about positive primes) so m-n=2 m=2+n m=2+(2k+1) m=2k+3 2j+1=2k+3 2j=2k+2 j=k+1 so we know j=k+1 so we can write our prime numbers as n=2k+1 m=2(k+1)+1=2k+2+1=2k+3 all the pairs will be of this form: (2k+3,2k+1)

myininaya (myininaya):

and i suppose we should assume k>0 since if k=0 we get 1 is in our set and 1 is not prime

myininaya (myininaya):

some of examples: so if k=1: (5,3) if k=2: (7,5) k=3: 3 will not work (we get a number that is not prime) k=4: 4 will not work k=5: 5 will not work k=6: 6 will not work k=7: 7 will not work k=8: (19,17) k=9: 9 will not work k=10: 10 will not work

myininaya (myininaya):

14 is the next one that works (31,29)

myininaya (myininaya):

then 20: (43,41)

myininaya (myininaya):

25: (53,51)

myininaya (myininaya):

28: (59,57) 29: (61,59) 35: (73,71) 43: (89,87) 44: (91,89) 50: (103,101)

myininaya (myininaya):

i think all these pairs are composed of primes

hero (hero):

You must have missed the part where it said the numbers had to be less than 50

myininaya (myininaya):

oops i did for k=50 lol

myininaya (myininaya):

yes you are right i forgot basically

hero (hero):

I finally found a mistake with you, computer

myininaya (myininaya):

lol

hero (hero):

This is the best day of my life

myininaya (myininaya):

you have been waiting for me to fail i see :(

hero (hero):

No, I haven't, lol....I love you

hero (hero):

You're my favorite computer

myininaya (myininaya):

:)

hero (hero):

How come you're always on here all the time? You could be working as a private tutor and making a lot of money....

myininaya (myininaya):

thanks for making the process quicker hero

myininaya (myininaya):

i don't like in person

myininaya (myininaya):

i like to choose if i want to help someone

hero (hero):

You can choose your own hours, etc...

myininaya (myininaya):

i don't wanna

myininaya (myininaya):

lol

hero (hero):

I'm not kidding